Consider a hydrogen atom modeled as a stationary p
Answer #1

Only force on electron is electrostatic force which is acting towards the center of the circle.

\frac{e^2}{4\pi \epsilon_0 r^2}=\frac{mv^2}{r}

\frac{e^2}{4\pi \epsilon_0 r}={mv^2} ---------------(1)

Energy of the hydrogen atom is E=PE+KE

E=\frac{-e^2}{4\pi \epsilon_0r}+\frac{mv^2}{2} ---------------------(2)

using (1) and (2)

E=\frac{-e^2}{4\pi \epsilon_0r}+\frac{e^2}{8\pi \epsilon_0r}=\frac{-e^2}{8\pi \epsilon_0r}

Energy required to ionize the hydrogen atom is e2 8.99×109(1.6×10-19)2 2x 5.291 77x10-11_ = 2.174 × 10-18 J